OpenStudy (anonymous):

∫e^(2t)sin(3t)dx, integral, can someone help me and explain with details!

OpenStudy (anonymous):

lol...is it dx or dt?

OpenStudy (anonymous):

sorry dt

OpenStudy (anonymous):

USe Integration by parts...

OpenStudy (anonymous):

could you explain more cuz it confused me

OpenStudy (anonymous):

Integration by parts... let u = sin3t dv = e^(2t) dt du = 3cos3t dt v = e^(2t)/2

OpenStudy (anonymous):

you should use intagration by parts: u=e^2t dv=sin3t dt & du=2e^2t dt v=-1/3cos(3t) -> uv-int(vdu) (e^2t)(-1/3cos(3t))+int((2/3)cos(3t)(e^2t))dt againu should use integration y parts u=e^2t dv=cos3t) &du=2e^2t v=1/3sin(3t) (e^2t)(1/3sin(3t))-int((1/3sin3t)(2e^2t))dt and we must name the phrase:((sin3t)(e^2t)) for example I and the equation solved